# Graph Condensation for Open-World Graph Learning

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Despite the potential to accelerate model training, current GC methods still fall short in their real-world practicality. By default, conventional GC methods are subsumed under a static setting, which requires that a large graph is firstly condensed to facilitate GNN training on the small graph, and the trained GNNs are then deployed on the same large graph for testing. Unfortunately, assuming the large graph stays unchanged throughout the entire process severely contradicts the dynamic and evolving nature of graph-structured data in the real world. In fact, graphs in many high throughput applications are inherently open-world (Geng et al., 2020; Bendale and Boult, 2016; Wu et al., 2020; Feng et al., 2023), where new nodes and classes continuously emerge and are integrated into the existing graph structure, as depicted in Figure 1. This phenomenon is exemplified in citation networks (Hu et al., 2020; Page et al., 1999), where some new papers explore established topics, while others venture into emerging areas. These papers in the novel areas are often developed from previous studies and cite a range of related literature. Such addition of nodes often introduce novel patterns that are distinctly different from those previously observed, increasing the graph’s complexity and diversity. Consequently, in addition to preserving performance on the initial large graph, GNNs trained on condensed graphs are expected to exhibit adaptability to novel patterns that emerge within dynamic environments. For instance, GC for neural architecture search (Zhou et al., 2022; Ding et al., 2022) initially condense a snapshot of the original graph, thereafter employing the condensed graph to accelerate the searching procedure and identify optimal model architecture. However, the evolving nature of graphs leads to a discrepancy between the model deployment environment and the condensed graph snapshot. Therefore, optimal model architecture identified based on condensed graphs should sustain its superior performance over time, even as the graph evolves. In a nutshell, the critical problem that arises for GC in open-world scenarios is: “How can GC methods be adapted to handle the dynamic nature of evolving graphs, ensuring that GNNs trained on condensed graphs remain accurate and robust in the face of continual changes in graph structures?
@@
In light of these challenges, we propose a novel graph condensation approach, Open-wold graph condensation (OpenGC), facilitating trained downstream GNNs to handle evolving graphs in the open-world scenarios. To tackle the distribution shift issue, we propose temporal invariance condensation that incorporates invariant learning (Arjovsky et al., 2019) to preserve the invariant patterns across different temporal environments in the condensed graph. To this end, a temporal data augmentation technique is designed to simulate the graph’s evolving pattern and exploit the structure-aware distribution shift by referring to the historic graph. It specifically considers the similarity of node pairs and targets the low-degree nodes, which are more susceptible to the influence of newly added neighbors. By this means, OpenGC can significantly enhance the adaptability of downstream GNNs by training them on the temporally generalized condensed graph, thus eliminating the need to laboriously design specific generalization modules for each GNN. In response to the existing sophisticated condensation procedure, we design an efficient GC paradigm that combines Kernel Ridge Regression (KRR) with non-parametric graph convolution, circumventing the nested loop optimization and heavy graph kernel-based encoding (Xu et al., 2023; Wang et al., 2024) in other KRR methods. Consequently, OpenGC is well-suited for managing life-long graph data and handling the continuous growth and dynamic changes of open-world graphs. The main contributions of this paper are threefold:
@@ 3.2. Structure-aware Environment Generation
where ${\mathbf{H}_{t,i}}$ and $\mathbf{H}_{t-1,i}$ are embeddings of node $i$ at task $T_{t}$ and $T_{t-1}$, respectively. As the graph evolves, the neighbors of node $i$ are extended, and its embedding changes from $\mathbf{H}_{t-1,i}$ to ${\mathbf{H}_{t,i}}$ after graph convolution. Therefore, $\Delta\mathbf{H}_{t,i}$ formulates the added neighbors in convolution and indicates the structure-aware distribution shift in task $T_{t}$. Then, we randomly select node $j$, which belongs to the same class as node $i$, and use its residual $\Delta\mathbf{H}_{t,j}$ to modify the embedding $\mathbf{H}_{t,i}$. The motivation hinges on the assumption that nodes in the same class follow a similar distribution shift pattern [^2] We use the drop edge and drop feature to construct the environments if no historic classes are available..
@@
where $\Delta\bar{\mathbf{H}}_{t,j}$ is the normalized residual and $\varepsilon_{t,i}$ controls the magnitude.
To determine the value of $\varepsilon_{t,i}$, we take into account the characteristics of the target nodes $i$ and the node $j$ responsible for generating the residual. Firstly, a higher weight is assigned to the target node $i$ if it has a lower degree, with the premise that nodes with fewer connections are more prone to being influenced, leading to a pronounced distribution shift upon the addition of new neighbor nodes. Moreover, the similarity between node $i$ and $j$ is assessed to promote a more significant distribution shift among similar nodes. Consequently, the magnitude is constructed as:
@@ 4. Experiments
We design comprehensive experiments to validate the efficacy of our proposed OpenGC and explore the following research questions: Q1: Compared to the SOTA GC methods, can OpenGC achieve better temporal generalization performance? Q2: Can the OpenGC condense the graph faster than other GC approaches? Q3: How do the different components, i.e., IRM and different constraints used in environments generation affect OpenGC? Q4: Can the condensed graph generated by OpenGC generalize well to different GNN architectures? Q5: How do the different hyper-parameters affect the OpenGC? Q6: What does the visualization of condensed nodes generated by OpenGC look like?
@@ 4.3. Condensation Time Comparison (Q2)
Figure 4. The hyper-parameter sensitivity analysis.
Figure 5. The visualization of t-SNE on condensed graph by OpenGC.
4.4. Ablation Study (Q3)

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671917isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Neural networks
@@ 1. Introduction
Despite the potential to accelerate model training, current GC methods still fall short in their real-world practicality. By default, conventional GC methods are subsumed under a static setting, which requires that a large graph is firstly condensed to facilitate GNN training on the small graph, and the trained GNNs are then deployed on the same large graph for testing. Unfortunately, assuming the large graph stays unchanged throughout the entire process severely contradicts the dynamic and evolving nature of graph-structured data in the real world. In fact, graphs in many high throughput applications are inherently open-world (Geng et al., 2020; Bendale and Boult, 2016; Wu et al., 2020; Feng et al., 2023), where new nodes and classes continuously emerge and are integrated into the existing graph structure, as depicted in Figure 1. This phenomenon is exemplified in citation networks (Hu et al., 2020; Page et al., 1999), where some new papers explore established topics, while others venture into emerging areas. These papers in the novel areas are often developed from previous studies and cite a range of related literature. Such addition of nodes often introduces novel patterns that are distinctly different from those previously observed, increasing the graph’s complexity and diversity. Consequently, in addition to preserving performance on the initial large graph, GNNs trained on condensed graphs are expected to exhibit adaptability to novel patterns that emerge within dynamic environments. For instance, GC for neural architecture search (Zhou et al., 2022; Ding et al., 2022) initially condenses a snapshot of the original graph, thereafter employing the condensed graph to accelerate the searching procedure and identify optimal model architecture. However, the evolving nature of graphs leads to a discrepancy between the model deployment environment and the condensed graph snapshot. Therefore, optimal model architecture identified based on condensed graphs should sustain its superior performance over time, even as the graph evolves. In a nutshell, the critical problem that arises for GC in open-world scenarios is: “How can GC methods be adapted to handle the dynamic nature of evolving graphs, ensuring that GNNs trained on condensed graphs remain accurate and robust in the face of continual changes in graph structures?
@@
In light of these challenges, we propose a novel graph condensation approach, Open-wold graph condensation (OpenGC), facilitating trained downstream GNNs to handle evolving graphs in the open-world scenarios. To tackle the distribution shift issue, we propose temporal invariance condensation that incorporates invariant learning (Arjovsky et al., 2019) to preserve the invariant patterns across different temporal environments in the condensed graph. To this end, a temporal data augmentation technique is designed to simulate the graph’s evolving pattern and exploit the structure-aware distribution shift by referring to the historic graph. It specifically considers the similarity of node pairs and targets the low-degree nodes, which are more susceptible to the influence of newly added neighbors. By this means, OpenGC can significantly enhance the adaptability of downstream GNNs by training them on the temporally generalized condensed graph, thus eliminating the need to laboriously design specific generalization modules for each GNN. In response to the existing sophisticated condensation procedure, we design an efficient GC paradigm that combines Kernel Ridge Regression (KRR) with non-parametric graph convolution, circumventing the nested loop optimization and heavy graph kernel encoding in other KRR-based GC methods (Xu et al., 2023; Wang et al., 2024). Consequently, OpenGC is well-suited for managing life-long graph data and handling the continuous growth and dynamic changes of open-world graphs. The main contributions of this paper are threefold:
@@ 3.2. Structure-aware Environment Generation
where ${\mathbf{H}_{t,i}}$ and $\mathbf{H}_{t-1,i}$ are embeddings of node $i$ at task $T_{t}$ and $T_{t-1}$, respectively. As the graph evolves, the neighbors of node $i$ are extended, and the node embedding changes from $\mathbf{H}_{t-1,i}$ to ${\mathbf{H}_{t,i}}$ after graph convolution. Therefore, $\Delta\mathbf{H}_{t,i}$ formulates the added neighbors in convolution and indicates the structure-aware distribution shift in task $T_{t}$, which can be leveraged to augment other node embeddings. Then, we randomly select node $j$, which belongs to the same class as node $i$, and use its residual $\Delta\mathbf{H}_{t,j}$ to modify the embedding $\mathbf{H}_{t,i}$. The motivation hinges on the assumption that nodes in the same class follow a similar distribution shift pattern [^2] We use the drop edge and drop feature to construct the environments if no historic classes are available..
@@
where $\Delta\bar{\mathbf{H}}_{t,j}$ is the normalized residual and $\varepsilon_{t,i}$ controls the magnitude. To determine the value of $\varepsilon_{t,i}$, we take into account the characteristics of the target nodes $i$ and the node $j$ responsible for generating the residual. Firstly, a higher weight is assigned to the target node $i$ if it has a lower degree, with the premise that nodes with fewer connections are more prone to being influenced, leading to a pronounced distribution shift upon the addition of new neighbor nodes. Moreover, the similarity between node $i$ and $j$ is assessed to promote a more significant distribution shift among similar nodes. Consequently, the magnitude is constructed as:
@@ 4. Experiments
We design comprehensive experiments to validate the efficacy of our proposed OpenGC and explore the following research questions: Q1: Compared to the SOTA GC methods, can OpenGC achieve better temporal generalization performance? Q2: Can the OpenGC condense the graph faster than other GC approaches? Q3: How do the different components, i.e., IRM and different constraints used in environment generation affect OpenGC? Q4: Can the condensed graph generated by OpenGC generalize well to different GNN architectures? Q5: How do the different hyper-parameters affect the OpenGC? Q6: What are the characteristics of the visualization of condensed nodes?
@@ 4.3. Condensation Time Comparison (Q2)
4.4. Ablation Study (Q3)
Figure 4. The visualization of t-SNE on condensed graph by OpenGC.